OpenStudy (loser66):

Find the mass of a cylinder \(0\leq r\leq 5,0\leq z\leq 10\), if its density at (x,y,z) is \(\sqrt{x^2+y^2}\) Please, help

OpenStudy (loser66):

@IrishBoy123

OpenStudy (irishboy123):

for small volume element in cylindrical \(dV = r dr d \theta dz\) mass of this is vol * density, so \(dm = \rho dV\) \(= \rho r dr d \theta dz\) and \(rho = r\) so \(= r^2 dr d \theta dz\) do you agree?

OpenStudy (irishboy123):

ie integrate that

OpenStudy (loser66):

Yup

OpenStudy (loser66):

ok, so, \(0\leq \theta\leq 2\pi\) \(0\leq r\leq 5\) \(0\leq z\leq10\) right?

OpenStudy (irishboy123):

yes

OpenStudy (loser66):

nope

OpenStudy (loser66):

the function should be r^(3/2) drdthetadz, right?

OpenStudy (loser66):

because rho at (x,y,z) = sqrt (x^2+y^2)= r my bad :(

OpenStudy (irishboy123):

i'm gonna type it up \(\int_{\theta = 0}^{ 2 \pi} \int_{z = 0}^{10} \int_{r = 0}^{5} r^2 dr dz d \theta\) https://www.wolframalpha.com/input/?i=int_%7Btheta+%3D+0%7D%5E%7B+2+pi%7D+int_%7Bz+%3D+0%7D%5E%7B10%7D+int_%7Br+%3D+0%7D%5E%7B5%7D++r%5E2+dr++dz+d+theta

OpenStudy (loser66):

yes, it is. omg. you are so so so good. How can you remember all of them??

OpenStudy (irishboy123):

lol!!

OpenStudy (irishboy123):

i suspect you know this but indulge me!!! latexing again

OpenStudy (irishboy123):

\(\int_{\theta = 0}^{ 2 \pi} \int_{z = 0}^{10} \int_{r = 0}^{5} r^2 dr dz d \theta\) \( = \int_{\theta = 0}^{ 2 \pi} d \theta * \int_{z = 0}^{10} dz * \int_{r = 0}^{5} r^2 dr \) so you can do these in your head!! how cool is that!!

OpenStudy (loser66):

yes, it is so simple now. :) Thanks a ton, friend.